OpenStudy (anonymous):

for each eqaution, identify the conic, and find the center, vertices, foci, and axis of symmetry, as appropriate. then, graph the equation. 9x^2+16y^2+36x-32y-92=0 [ALGEBRA 2 HELP] MEDALS GIVEN.

OpenStudy (anonymous):

8x^2-3y^2=48 x^2-2x-2y-1=0 2x^2+2y^2-3x-5y+3=0